OpenStudy (anonymous):

Solve the following difference equations (show your solutions): a. an+1=ran, a0=b b. an+1=ran+c, a0=b

OpenStudy (anonymous):

\[a _{o}\]

OpenStudy (anonymous):

confused. is it \(a_{n+1}=ra_n\) a geometric sequence?

OpenStudy (anonymous):

general term looks like \(a_n=r^na_0\) or i guess \(r^nb\) since \(a_0=b\)